HS-H BOND DISSOCIATION ENERGY DETERMINATION BY THRESHOLD ION-PAIR PRODUCTION SPECTROSCOPY (TIPPS)

\underline{X. K. HU}, R. C. SHIELL, Q. J. HU and J. W. HEPBURN, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

The first single photon threshlod ion-pair production spectrum (TIPPS) of a polyatomic molecule, H_{2}S, was recorded by monitoring H^{+} ion generated from the dissociation process H_{2}S → HS^{−} + H^{+}. Its ion-pair threshold was determined to be IPP = 122458 ± 3 cm^{−1}. As a result, the bond dissociation energy (D(HS-H) = 31451 ± 4 cm^{−1}) could be calcula ted from D(HS-H) = IPP - IP(H) + EA(HS), by knowing the ionization potential of the hydrogen atom (IP(H) = 109678.772 cm^{−1}) and the electron affinity of the HS radical (EA(HS) = 18672 ± 2 cm^{−1}).